OpenStudy (anonymous):

After 1 second, the missile is 224 feet in the air; after 2 seconds, it is 416 feet in the air. using the quadratic function, find the height in feet, of the missile after 10 seconds in the air. missile was launched from the ground

OpenStudy (anonymous):

Assume: height = at^2 + bt + c We have three sets of data to use: t = 1, h = 224 t = 2, h = 416. AND an important one not to miss, t = 0, h = 0 So what can we do? Well, using this last one, we can clearly see 0 = 0a + 0b + c, so c = 0 => height = at^2 + bt Now, using the previous two, we can create two more equations: t=1, h = 224 => 224 = a + b (*) t =2,h = 416 => 416 = 4a + 2b (**) What can we do now? SOLVE equations (*) and (**) for a and b!!! Then using these values in the general equation with t = 10, find the height then!!!